OpenStudy (coconut):

What is the sum of the arithmetic sequence 8, 14, 20..., if there are 22 terms?

OpenStudy (anonymous):

a1 = 8 a22 = 8 + 6*21 a22 = 8 + 126 a22 = 134 formula : (n+1 ) * (a0 + an) /2 a0 = 8 an = 134 you can do the rest iguess :)

OpenStudy (anonymous):

I would solve like this: find a formula for the sequence. an=n*6+2 seems to do. Then, sum from i = 1 to 22: (i*6 + 2). Manipulate that sum a bit. Use some properties of summation to break it down. You should reach 44+6 * (1+2+3+....+22). There's a formula for the last bit.

OpenStudy (coconut):

my options are 1,558 1,560 1,562 1,564

OpenStudy (anonymous):

better get calculating then.

OpenStudy (anonymous):

for arithmetic sequences, an = a1+ (n-1)d and Sn= (a1+an)(n)/2 d=6, a22=8+21(6)=134 S22=(8+134)(22)/2= 142(11)= 1562

OpenStudy (anonymous):

so it was 1562